You are concerned about the possibility of experiencing on earthquake during the next year. The probability of an earthquake is 0.5%. Your local insurer offers to pay you USD300, 000 if your house is damaged as a result of an earthquake. The insurance policy costs USD1, 500.00. If the inflation rate is forecast to be 10% during the next year, is the price of insurance policy fair.
Answer #1

Payoff if earthquake happens next year=$300000 and associated probability is 0.5%

Payoff if earthquake does not happens next year=0 and associated probability is (100%-0.5%)=99.5%

Expected payoff for next year= 0.5%*300000+99.5%*0=$1500

Present value of expected payoff =Future value/(1+inflation rate)=1500/(1+10%)=$1363.63

However, the policy costs now is $1500 in comparison with the present value of expected payoff of $1363.63

Hence, insurance policy is overpriced and it's not fair.